# Production of dileptonic bound states in the Higgs boson decay

### v1

@@ II Production of single leptonic bound states
Figure 1: Different mechanisms of single dilepton production in the Higgs boson decay.
The lepton-photon (lepton-Z-boson) mechanism of single production of dimuonium and ditauonium is determined by three amplitudes presented in Fig. 1(a,b,c). The initial expression for the amplitude of such a process can be represented as a convolution of the amplitude of the production of a photon and a lepton-antilepton pair with the Bethe-Salpeter wave function, which describes, within the framework of the four-dimensional formalism, a lepton-antilepton bound state moving with 4-momentum $P$. Transforming the amplitude during the transition to three-dimensional formalism and using the law of transformation of the quasipotential wave function during the transition from the moving frame of reference to the rest frame of the bound state of leptons, we obtain two decay amplitudes (Fig. 1(a,b)) in the form:
${\cal M}_{1}(k,P)=\frac{4\pi\alpha m_{l}}{\sin 2\theta_{W}M_{Z}}\int\frac{d{\bf p}}{(2\pi)^{3}}\frac{\Psi({\bf p})}{\frac{\varepsilon}{m}\frac{(\varepsilon+m)}{2m}[(r-p_{2})^{2}-m^{2}]}\times$
$Tr\bigl{\{}[\frac{\hat{v}-1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}-\frac{\hat{p}}{2m}]\hat{\varepsilon}_{l\bar{l}}\frac{(\hat{v}+1)}{2\sqrt{2}}[\frac{\hat{v}+1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}+\frac{\hat{p}}{2m}]\hat{\varepsilon}_{\gamma}(\hat{r}-\hat{p}_{2}+m)\bigr{\}},$ (2)
${\cal M}_{2}(k,P)=\frac{4\pi\alpha m_{l}}{\sin 2\theta_{W}M_{Z}}\int\frac{d{\bf p}}{(2\pi)^{3}}\frac{\Psi({\bf p})}{\frac{\varepsilon}{m}\frac{(\varepsilon+m)}{2m}[(p_{1}-r)^{2}-m^{2}]}\times$
$Tr\bigl{\{}[\frac{\hat{v}-1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}-\frac{\hat{p}}{2m}]\hat{\varepsilon}_{l\bar{l}}\frac{(\hat{v}+1)}{2\sqrt{2}}[\frac{\hat{v}+1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}+\frac{\hat{p}}{2m}](\hat{p}_{1}-\hat{r}+m)\hat{\varepsilon}_{\gamma}\bigr{\}},$ (3)
where $r$ is the Higgs boson four-momentum, $\varepsilon_{\gamma}$ is the photon polarization vector, $\varepsilon_{l\bar{l}}$ is the polarization vector of dimuonium (ditauonium), $k$ is the photon four-momentum, $\varepsilon(p)=\sqrt{p^{2}+m^{2}}$, $v=P/M$, $M$ is the mass of $l\bar{l}$ state. $p$ is the lepton relative four-momentum. The expressions for the amplitudes (II), (II) are obtained as a result of the Lorentz transformations in the same way as in our works [13, 14, 15]. They make it possible to take into account relativistic corrections connected with the momentum of relative motion of leptons $p$.
The simplifications that can be made in the amplitudes are primarily related to the denominators of the lepton propagators. The magnitude of the relative momentum is small compared to the mass of the Higgs boson $M_{H}$, so we have
@@
where the lepton mass can be approximately replaced here by $m=\frac{M}{2}$ neglecting bound state effects.
@@
where for convenience the particle mass ratios are introduced: $r_{1}=\frac{m}{M}$, $r_{2}=\frac{M_{H}}{M}$.
Relativistic corrections in (II) of the second and fourth orders in $p/m$ are determined in our approach by the parameters $\omega_{1}$, $\omega_{2}$ and the value of the wave function at zero $\tilde{\psi}(0)$. The definition of these parameters and their numerical calculation is given in the next section. Note that we have retained the 4th order $\alpha$ corrections here only to demonstrate the overall amplitude structure. In what follows, when obtaining numerical estimates, we neglect fourth-order corrections.
Using the decay amplitude (II), we can calculate the differential decay width according to the formula:
@@
Three-momentum of final dilepton state is expressed in terms of masses of higgs boson $(M_{H})$, $Z$-boson $(M_{Z})$, and dilepton state $(M)$:
@@
In the numerator of the amplitude (II) we define two functions $g_{1,2}$ as follows:
@@
After summing over the polarizations of the photon and ortho-dileptonium we obtain total decay width $H\to\gamma+(l\bar{l})$ in the form:
@@
To obtain total width of the decay $H\to Z+(l\bar{l})$, including the contribution of the $ZZ$-production mechanism (see, Fig. 1(c)), it is necessary to make the following substitution in (II):
@@
The expressions (8), (9)-(II) clearly show that the main parameters that determine numerical values of decay branchings include, together with the fine structure constant, the particle mass ratios $r_{1}$, $r_{2}$, $r_{3}$.
Figure 2: Quark loop and $W$-boson loop mechanisms of single dilepton bound state production in the Higgs boson decay.
The studies we carried out in [13, 14] show that it is necessary to take into account the loop ($W$-boson and quark) production mechanism $Z+(l\bar{l})$ shown in Fig. 4. Although formally an additional degree of $\alpha$ appears in it compared to the $ZZ$-boson mechanism, the fine structure constant is not a decisive factor in determining the order of magnitude of the contribution. Numerous particle mass ratio parameters begin to play a more significant role here. The loops of heavy quarks and the $W$-boson make their contribution to the structure functions $g_{1}$ and $g_{2}$, which can be represented as follows:
@@
$\Gamma=\frac{8\pi^{3}\alpha^{4}(r_{2}^{2}-1)ctg^{2}\theta_{W}M_{Z}^{2}}{r_{2}^{3}M^{4}}|\tilde{\psi}(0)|^{2}A_{W}^{2}\bigl{[}3|g_{1W}|^{2}-\frac{1}{4}(r_{2}^{2}-1)^{2}|g_{2W}|^{2}\bigr{]},$ (20)
@@ III Pair production of leptonic bound states
In addition to the processes of single production of bound states of leptons, we also consider processes of pair production of orto-positronium ($e^{+}e^{-}$), orto-dimuonium ($\mu^{+}\mu^{-}$), orto-ditauonium ($\tau^{+}\tau^{-}$) in the Higgs boson decay. There are five different production mechanisms presented in Figs. 3, 4, 5, 6, 7, which we consider below.
@@ III.1 Lepton - photon production mechanism A
Figure 3: Lepton - photon mechanism A of pair dilepton production in Higgs boson decay.
Taking into account the law of transformation of the wave function of a bound state upon transition to a rest frame, we obtain wave functions in (III.1) in the form [13, 14]:
@@
where we introduce two auxiliary 4-vectors $v_{1}=\frac{P}{M}$, $v_{2}=\frac{Q}{M}$, $\varepsilon(p)=\sqrt{{\bf p}^{2}+m^{2}}$, $m$ is the lepton mass. In equations (III.1), (III.1) the projection operators on $(^{3}S_{1})$ dilepton states are introduced. Projection operators are constructed in terms of the Dirac bispinors in the rest frame as follows:
$\Pi_{S=1}=\frac{(1+\hat{v}_{1})}{2\sqrt{2}}{\hat{\varepsilon}}_{1}(P,S_{z}).$ (29)
The denominators of propagators in amplitude (III.1) can be simplified due to the condition $M_{h}\gg m$ as follows:
@@
The calculation of trace in (III.1) of the product of $\gamma$-factors and various simplifications are performed in the Form package [21]. Considering small factors $\frac{{\bf p}}{m}$, $\frac{{\bf q}}{m}$, we perform an expansion on them in the numerator of the amplitude, while preserving the second-order terms. Then the numerator of this amplitude takes the form:
@@
Special relativistic parameters $\omega_{1}$, $\tilde{\psi}(0)$ take into account relativistic corrections connected with the relative motion of leptons. The calculation of such parameters is discussed in details below in section IV.1. After all transformations, the production amplitude (III.1) takes the form:
@@ III.2 Lepton - Z-boson production mechanism B
Figure 4: Lepton - $Z$-boson mechanism B of pair dilepton production in the Higgs boson decay.
@@ III.3 Lepton - photon production mechanism C
Figure 5: Lepton - photon mechanism C of pair dilepton production in the Higgs boson decay.
@@ III.4 Lepton - Z-boson production mechanism D
Figure 6: Lepton - $Z$-boson mechanism D of pair dilepton production in the Higgs boson decay.
@@
$\frac{1}{k_{4}^{2}-M_{Z}^{2}}=\frac{1}{(q_{1}+q_{2})^{2}-M_{Z}^{2}}\approx\frac{1}{M^{2}-M_{Z}^{2}},~{}~{}\frac{1}{(-r+p_{1})^{2}-m^{2}}\approx\frac{1}{(r-p_{2})^{2}-m^{2}}\approx\frac{2}{M_{H}^{2}}.$ (47)
After calculating two traces in the numerator (III.4) we obtain:
@@
Then the production amplitude (III.4) has the same general structure as previous amplitudes:
@@ III.5 ZZ-boson production mechanism E
In this case, we have direct (a) and crossed (b) amplitudes (see Fig.7):
@@
Figure 7: Pair dilepton ZZ-boson production mechanism in the Higgs boson decay.
@@ III.6 Quark loop and W-boson loop production mechanism
Figure 8: Pair dilepton production mechanism in the Higgs boson decay with quark and $W$-boson loop.
@@
where the general factor corresponds to amplitude (III.6) so that the functions $g^{W}_{1}$ and $g^{W}_{2}$ are included in the complete functions $g_{1}^{tot}$ and $g_{2}^{tot}$ with unit coefficients.
@@ IV Relativistic corrections
Relativistic corrections represent an important class of corrections for the production of bound quark states. In the case of the production of bound states of leptons, such corrections are small, since they are determined by the value of the fine structure constant. Taking them into account is connected with increasing the accuracy of calculating the decay width.
When calculating decay widths, we take into account relativistic corrections of several types. The first group of relativistic corrections is connected with the law of transformation of the wave function of a bound state into a rest frame and the momenta of the relative motion of leptons in the decay amplitude. The second group is determined by the Breit potential.
@@ IV.2 Relativistic corrections to the wave function
where ${\tilde{G}}_{1S}$ is reduced Coulomb Green function, $\psi^{(0)}_{1S}$ is defined by (65), $\Delta V$ is perturbative potential.
@@
where $x=Wr$.
The Breit potential contains several terms, the contribution of which must be calculated:
@@
It is convenient to start calculating corrections to the wave function at zero with the $\delta$ term of the potential:
@@
$\tilde{G}_{1S}({\bf p},{\bf q})=-\frac{64\pi}{\alpha W^{4}}\left[\frac{\pi^{2}W^{5}\delta({\bf p}-{\bf q})}{4({\bf p}^{2}+W^{2})}+\frac{W^{6}}{4({\bf p}^{2}+W^{2})({\bf p}-{\bf q})^{2}({\bf q}^{2}+W^{2})}+R({\bf p},{\bf q})\right]~{}~{}~{}.$ (78)
In parallel with this correction, let us also consider the contribution from the second term (73):
$\psi^{(1)}_{2}(0)=-\frac{1}{4m^{3}}\int\tilde{G}_{1S}(0,{\bf r})\hat{\bf p}^{4}\psi(r)d{\bf r})=~{}$
$-\frac{\pi\alpha}{m^{2}}\psi^{C}(0)\int\tilde{G}({\bf q},{\bf p})\frac{d{\bf p}d{\bf q}}{(2\pi)^{6}}\left[1-\frac{W^{2}(2p^{2}+W^{2})}{(p^{2}+W^{2})^{2}}\right].$ (79)
From expressions (77) , (IV.2) it follows that the divergent terms in the matrix elements cancel each other so that the contribution of the operator from (73) will be finite. After calculating three integrals with functions from (78) , we obtain the following summary result:
$\Delta\psi^{(1)}(0)=\psi^{C}(0)[-\frac{63}{128}\alpha^{2}].$ (80)
@@
$=\psi^{C}(0)\frac{-2\mu\alpha^{2}}{\pi m^{2}}\int\tilde{G}({\bf q},{\bf p})\frac{d{\bf q}}{(2\pi)^{3}}\frac{d{\bf p}}{(2\pi)^{3}}\frac{d{\bf p^{\prime}}}{({\bf p}-{\bf p^{\prime}})^{2}}\left[\frac{1}{(p^{\prime 2}+W^{2})}-\frac{W^{2}}{(p^{\prime 2}+W^{2})^{2}}\right].~{}~{}~{}$
@@
$\Delta\psi^{(2)}(0)=\psi^{C}(0)\frac{4W^{2}}{\pi m^{2}}\int^{\sim m}\frac{d{\bf q}d{\bf p}}{(2\pi)^{3}(p^{2}+W^{2})({\bf p}-{\bf p^{\prime}})^{2}(p^{\prime 2}+W^{2})}=\psi^{C}(0)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}.$ (83)
@@
Calculating the momentum integral, we obtain in leading logarithmic order:
$\Delta\psi_{rel}(0)=\psi^{C}(0)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}.$ (85)
@@
$\Delta\psi_{rel}^{(1)}=\psi^{C}(0)\left[-\frac{\alpha^{2}}{8}\right],~{}~{}~{}$ (86)
$\Delta\psi_{rel}^{(2)}=\psi^{C}(0)\frac{W^{2}}{2m^{2}\pi^{6}}\int\frac{p^{\prime 2}d{\bf p^{\prime}}d{\bf p}d{\bf q}}{({\bf p}-{\bf p^{\prime}})^{2}(p^{\prime 2}+1)^{2}(q^{2}+1)({\bf q}-{\bf p})^{2}(p^{2}+1)}=$ (87)
$\psi^{C}(0)\left[\frac{\alpha^{2}}{12}(\pi^{2}-\frac{3}{2})\right],$
@@
$\Delta\psi_{2}=\psi^{C}(0)\left[\frac{1}{2}\alpha^{2}\ln\alpha^{-1}+\frac{5}{8}\alpha^{2}\right].$ (89)
@@
$\Delta\psi^{(3)}(0)=\psi^{C}(0)\left[2-\frac{7}{3}S(S+1)\right]\frac{1}{4}\alpha^{2}\ln\alpha^{-1}.$ (90)
@@
$\psi(0)=\psi^{C}(0)\left[a_{vp}\frac{\alpha}{\pi}\right],~{}~{}~{}r_{6}=\frac{m_{e}}{m\alpha},$ (92)
@@
Summing contributions (86), (87), (88), (89), (90) and (93), we obtain a correction to the wave function at the zero of the bound state of leptons as follows:
$\psi(0)=\psi^{C}(0)\left\{1+a_{vp}\frac{\alpha}{\pi}+\left(2-\frac{7}{6}S(S+1)\right)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}-\frac{3}{128}\alpha^{2}\right\}.$ (94)

### v2

@@ II Production of single leptonic bound states
Figure 1: Different mechanisms of single leptonium production in the Higgs boson decay.
The lepton-photon (lepton-Z-boson) mechanism of single production of dimuonium and ditauonium is determined by three amplitudes presented in Fig. 1(a,b,c). The initial expression for the amplitude of such a process can be represented as a convolution of the amplitude of the production of a photon and a lepton-antilepton pair with the Bethe-Salpeter wave function, which describes, within the framework of the four-dimensional formalism, a lepton-antilepton bound state moving with 4-momentum $P$. Transforming the amplitude during the transition to three-dimensional formalism and using the law of transformation of the quasipotential wave function during the transition from the moving frame of reference to the rest frame of the lepton bound state (see, [17, 18]), we obtain two decay amplitudes (Fig. 1(a,b)) in the form:
${\cal M}_{1}(k,P)=\frac{4\pi\alpha m_{l}}{\sin 2\theta_{W}M_{Z}}\int\frac{d{\bf p}}{(2\pi)^{3}}\frac{\Psi({\bf p})}{\frac{\varepsilon}{m}\frac{(\varepsilon+m)}{2m}[(r-p_{2})^{2}-m^{2}]}\times$ (2)
$Tr\bigl{\{}[\frac{\hat{v}-1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}-\frac{\hat{p}}{2m}]\hat{\varepsilon}_{l\bar{l}}\frac{(\hat{v}+1)}{2\sqrt{2}}[\frac{\hat{v}+1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}+\frac{\hat{p}}{2m}]\hat{\varepsilon}_{\gamma}(\hat{r}-\hat{p}_{2}+m)\bigr{\}},$
${\cal M}_{2}(k,P)=\frac{4\pi\alpha m_{l}}{\sin 2\theta_{W}M_{Z}}\int\frac{d{\bf p}}{(2\pi)^{3}}\frac{\Psi({\bf p})}{\frac{\varepsilon}{m}\frac{(\varepsilon+m)}{2m}[(p_{1}-r)^{2}-m^{2}]}\times$ (3)
$Tr\bigl{\{}[\frac{\hat{v}-1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}-\frac{\hat{p}}{2m}]\hat{\varepsilon}_{l\bar{l}}\frac{(\hat{v}+1)}{2\sqrt{2}}[\frac{\hat{v}+1}{2}-\hat{v}\frac{p^{2}}{2m(\varepsilon+m)}+\frac{\hat{p}}{2m}](\hat{p}_{1}-\hat{r}+m)\hat{\varepsilon}_{\gamma}\bigr{\}},$
where $r$ is the Higgs boson four-momentum, $\varepsilon_{\gamma}$ is the photon polarization vector, $\varepsilon_{l\bar{l}}$ is the polarization vector of dimuonium (ditauonium), $k$ is the photon four-momentum, $\varepsilon(p)=\sqrt{p^{2}+m^{2}}$, $v=P/M$, $M$ is the mass of $l\bar{l}$ state. $p$ is the lepton relative four-momentum. The expressions for the amplitudes (2), (3) are obtained as a result of the Lorentz transformations in the same way as in our works [13, 14, 15]. They make it possible to take into account relativistic corrections connected with the momentum of relative motion of leptons $p$. The 4-momenta of the photon and leptonium (bound state lepton-anti-lepton) are given as arguments in the amplitudes (2), (3), although it should be remembered that they lie on the mass surface, and the following conservation law is satisfied: $r=k+P$.
The simplifications that can be made in the amplitudes are primarily related to the denominators of the lepton propagators. The magnitude of the relative momentum is small compared to the mass of the Higgs boson $M_{H}$, so we have:
@@
where the lepton mass can be approximately replaced here by $m=\frac{M}{2}$ neglecting bound state effects. We completely neglect corrections of the form $|{\bf p}|/M_{H}$. At the same time, we keep further in the decay amplitudes the second-order correction for small ratios $|{\bf p}|/m$ relative to the leading order result. If it is necessary to increase the accuracy of the calculation, the corrections $|{\bf p}|/M_{H}$ can also be taken into account within the used method.
@@
where for convenience the particle mass ratios are introduced: $r_{1}=\frac{m}{M}$, $r_{2}=\frac{M_{H}}{M}$. The expression (5) clearly demonstrates the general structure of the decay amplitudes, which contains two characteristic parts with $(v_{\gamma}\varepsilon)(v\varepsilon_{\gamma})$ and $(\varepsilon_{\gamma}\varepsilon_{l\bar{l}})$.
Relativistic corrections in (5) of the second and fourth orders in $p/m$ are determined in our approach by the parameters $\omega_{1}$, $\omega_{2}$ and the value of the wave function at zero $\tilde{\psi}(0)$. The definition of these parameters and their numerical calculation is given in the next section. Note that we have retained the 4th order $\alpha$ corrections here only to demonstrate the overall amplitude structure. In what follows, when obtaining numerical estimates, we neglect fourth-order corrections.
Using the decay amplitude (5), we can calculate the differential decay width according to the formula:
@@
Three-momentum of final leptonium state is expressed in terms of masses of the Higgs boson $(M_{H})$, $Z$-boson $(M_{Z})$, and $(l\bar{l})$ state $(M)$:
@@
In the numerator of the amplitude (5) we define two functions $g_{1,2}$ as follows:
@@
After summing over the polarization of the photon and ortho-dileptonium we obtain total decay width $H\to\gamma+(l\bar{l})$ in the form:
@@
To obtain total width of the decay $H\to Z+(l\bar{l})$, including the contribution of the $ZZ$-production mechanism (see, Fig. 1(c)), it is necessary to make the following substitution in (10):
@@
The expressions (8), (9)-(11) clearly show that the main parameters that determine numerical values of decay branchings include, together with the fine structure constant, the particle mass ratios $r_{1}$, $r_{2}$, $r_{3}$.
Figure 2: Quark loop and $W$-boson loop mechanisms of single leptonium production in the Higgs boson decay.
The studies we carried out in [13, 14] show that it is necessary to take into account the loop ($W$-boson and quark) production mechanism $Z+(l\bar{l})$ shown in Fig. 4. Although formally an additional degree of $\alpha$ appears in it compared to the $ZZ$-boson mechanism, the fine structure constant is not a decisive factor in determining the order of magnitude of the contribution. It is useful to mention here that a loop mechanism with W-boson and quark loops is needed to describe the decay process $H\to Z+\gamma$, the first evidence for which was presented recently in [19]. Numerous particle mass ratio parameters begin to play a more significant role in loop production mechanism. The loops of heavy quarks and the $W$-boson make their contribution to the structure functions $g_{1}$ and $g_{2}$, which can be represented as follows:
@@
$\Gamma_{W}=\frac{8\pi^{3}\alpha^{4}(r_{2}^{2}-1)ctg^{2}\theta_{W}M_{Z}^{2}}{r_{2}^{3}M^{4}}|\tilde{\psi}(0)|^{2}A_{W}^{2}\bigl{[}3|g_{1W}|^{2}-\frac{1}{4}(r_{2}^{2}-1)^{2}|g_{2W}|^{2}\bigr{]},$ (20)
@@ III Pair production of leptonic bound states
In addition to the processes of single production of bound states of leptons, we also consider processes of pair production of orto-positronium ($e^{+}e^{-}$), orto-dimuonium ($\mu^{+}\mu^{-}$), orto-ditauonium ($\tau^{+}\tau^{-}$) in the Higgs boson decay. There are five different production mechanisms presented in Figs. (3 - 7), which we consider below.
@@ III.1 Lepton - photon production mechanism A
Figure 3: Lepton - photon mechanism A of pair leptonium production in Higgs boson decay.
Taking into account the law of transformation of the wave function of a bound state upon transition to a rest frame, we obtain wave functions in (26) in the form [13, 14]:
@@
where we introduce two auxiliary 4-vectors $v_{1}=\frac{P}{M}$, $v_{2}=\frac{Q}{M}$, $\varepsilon(p)=\sqrt{{\bf p}^{2}+m^{2}}$, $m$ is the lepton mass. In equations (27), (28) the projection operators on $(^{3}S_{1})$ dilepton states are introduced. Projection operators are constructed in terms of the Dirac bispinors in the rest frame as follows:
$\Pi_{S=1}=\frac{(1+\hat{v}_{i})}{2\sqrt{2}}{\hat{\varepsilon}}_{i}(P,S_{z}).$ (29)
The denominators of propagators in amplitude (26) can be simplified due to the condition $M_{H}\gg m$ as follows:
@@
The calculation of trace in (26) of the product of $\gamma$-factors and various simplifications are performed in the Form package [24]. Considering small factors $\frac{{\bf p}}{m}$, $\frac{{\bf q}}{m}$, we perform an expansion on them in the numerator of the amplitude, while preserving the second-order terms. Then the numerator of this amplitude takes the form:
@@
Special relativistic parameters $\omega_{1}$, $\tilde{\psi}(0)$ take into account relativistic corrections connected with the relative motion of leptons. The calculation of such parameters is discussed in details below in section IV. After all transformations, the production amplitude (26) takes the form:
@@ III.2 Lepton - Z-boson production mechanism B
Figure 4: Lepton - $Z$-boson mechanism B of pair leptonium production in the Higgs boson decay.
@@ III.3 Lepton - photon production mechanism C
Figure 5: Lepton - photon mechanism C of pair leptonium production in the Higgs boson decay.
@@ III.4 Lepton - Z-boson production mechanism D
Figure 6: Lepton - $Z$-boson mechanism D of pair leptonium production in the Higgs boson decay.
@@
$\frac{1}{k_{4}^{2}-M_{Z}^{2}}=\frac{1}{(q_{1}+q_{2})^{2}-M_{Z}^{2}}\approx\frac{1}{M^{2}-M_{Z}^{2}},~{}~{}$
$\frac{1}{(-r+p_{1})^{2}-m^{2}}\approx\frac{1}{(r-p_{2})^{2}-m^{2}}\approx\frac{2}{M_{H}^{2}}.$ (47)
After calculating two traces in the numerator (46) we obtain:
@@
Then the production amplitude (46) has the same general structure as previous amplitudes:
@@ III.5 ZZ-boson production mechanism E
In this case, there are direct (a) and crossed (b) amplitudes (see Fig.7):
@@
Figure 7: Pair leptonium ZZ-boson production mechanism in the Higgs boson decay.
@@
The indices dir and cr denote the contributions of the direct and crossed interaction amplitudes in Fig.8.
@@ III.6 Quark loop and W-boson loop production mechanism
Figure 8: Pair leptonium production mechanism in the Higgs boson decay with quark and $W$-boson loop.
@@
where the general factor corresponds to amplitude (58) so that the functions $g^{W}_{1}$ and $g^{W}_{2}$ are included in the complete functions $g_{1}^{tot}$ and $g_{2}^{tot}$ with unit coefficients.
@@ IV Relativistic corrections
Relativistic corrections represent an important class of corrections for the production of bound quark states. In the case of the production of lepton bound states, such corrections are small, since they are determined by the value of the fine structure constant. Taking them into account is connected with increasing the accuracy of calculating the decay width.
When calculating decay widths, we take into account relativistic corrections of several types. The first group of relativistic corrections is connected with the law of transformation of the wave function of a bound state into a rest frame and the momenta of the relative motion of leptons in the decay amplitude. The second group is determined by the Breit potential in the wave function of the bound state in the rest frame.
@@ IV.2 Relativistic corrections to the wave function
where ${\tilde{G}}_{1S}$ is the reduced Coulomb Green’s function, $\psi^{(0)}_{1S}$ is defined by (65), $\Delta V$ is the perturbative potential.
@@
where $x=Wr$, $\gamma_{E}$ is the Euler constant.
The Breit potential contains several terms, the contribution of which must be calculated according to formula (70):
@@
It is convenient to start calculating corrections to the wave function at zero with the $\delta$ term of the potential $\Delta V_{1}$:
@@
$\tilde{G}_{1S}({\bf p},{\bf q})=-\frac{64\pi}{\alpha W^{4}}\left[\frac{\pi^{2}W^{5}\delta({\bf p}-{\bf q})}{4({\bf p}^{2}+W^{2})}+\frac{W^{6}}{4({\bf p}^{2}+W^{2})({\bf p}-{\bf q})^{2}({\bf q}^{2}+W^{2})}+R({\bf p},{\bf q})\right].$ (78)
In parallel with this correction, let us also consider the contribution from the second term of operator (73):
$\psi^{(1)}_{2}(0)=-\frac{1}{4m^{3}}\int\tilde{G}_{1S}(0,{\bf r})\hat{\bf p}^{4}\psi^{C}_{1S}(r)d{\bf r})=~{}$
$-\frac{\pi\alpha}{m^{2}}\psi^{C}_{1S}(0)\int\tilde{G}({\bf q},{\bf p})\frac{d{\bf p}d{\bf q}}{(2\pi)^{6}}\left[1-\frac{W^{2}(2p^{2}+W^{2})}{(p^{2}+W^{2})^{2}}\right].$ (79)
From expressions (77) , (79) it follows that the divergent terms in the matrix elements cancel each other so that the contribution of the operator from (73) will be finite. After calculating three integrals with functions from (78) , we obtain the following summary result:
$\Delta\psi^{(1)}(0)=\psi^{C}_{1S}(0)[-\frac{63}{128}\alpha^{2}].$ (80)
@@
$=\psi^{C}_{1S}(0)\frac{-2\mu\alpha^{2}}{\pi m^{2}}\int\tilde{G}({\bf q},{\bf p})\frac{d{\bf q}}{(2\pi)^{3}}\frac{d{\bf p}}{(2\pi)^{3}}\frac{d{\bf p^{\prime}}}{({\bf p}-{\bf p^{\prime}})^{2}}\left[\frac{1}{(p^{\prime 2}+W^{2})}-\frac{W^{2}}{(p^{\prime 2}+W^{2})^{2}}\right].~{}~{}~{}$
@@
$\Delta\psi^{(2)}(0)=\psi^{C}_{1S}(0)\frac{4W^{2}}{\pi m^{2}}\int^{\sim m}\frac{d{\bf q}d{\bf p}}{(2\pi)^{3}(p^{2}+W^{2})({\bf p}-{\bf p^{\prime}})^{2}(p^{\prime 2}+W^{2})}=\psi^{C}_{1S}(0)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}.$ (83)
@@
Calculating the momentum integral, we obtain in leading logarithmic order the result:
$\Delta\psi_{rel}(0)=\psi^{C}_{1S}(0)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}.$ (85)
@@
$\Delta\psi_{rel}^{(1)}=\psi^{C}_{1S}(0)\left[-\frac{\alpha^{2}}{8}\right],$ (86)
$\Delta\psi_{rel}^{(2)}=\psi^{C}_{1S}(0)\frac{W^{2}}{2m^{2}\pi^{6}}\int\frac{p^{\prime 2}d{\bf p^{\prime}}d{\bf p}d{\bf q}}{({\bf p}-{\bf p^{\prime}})^{2}(p^{\prime 2}+1)^{2}(q^{2}+1)({\bf q}-{\bf p})^{2}(p^{2}+1)}=$ (87)
$\psi^{C}_{1S}(0)\left[\frac{\alpha^{2}}{12}(\pi^{2}-\frac{3}{2})\right],$
@@
$\Delta\psi_{2}=\psi^{C}_{1S}(0)\left[\frac{1}{2}\alpha^{2}\ln\alpha^{-1}+\frac{5}{8}\alpha^{2}\right].$ (89)
@@
$\Delta\psi^{(3)}(0)=\psi^{C}_{1S}(0)\left[2-\frac{7}{3}S(S+1)\right]\frac{1}{4}\alpha^{2}\ln\alpha^{-1},$ (90)
where $S$ is the spin of the bound state which is equal to 1 for orto-leptonium.
@@
$\psi(0)=\psi^{C}_{1S}(0)\left[a_{vp}\frac{\alpha}{\pi}\right],~{}~{}~{}r_{6}=\frac{m_{e}}{m\alpha},$ (92)
@@
Summing contributions (86) - (90) and (93), we obtain a correction to the wave function at the zero of the bound state of leptons as follows:
$\psi(0)=\psi^{C}_{1S}(0)\left\{1+a_{vp}\frac{\alpha}{\pi}+\left(2-\frac{7}{6}S(S+1)\right)\frac{1}{2}\alpha^{2}\ln\alpha^{-1}-\frac{3}{128}\alpha^{2}\right\}.$ (94)